OpenStudy (hba):

A triangle is formed by the x-axis,the y-axis and the line y=ax+b.If a=-b ,then what is the volume of the cone generated by rotating this triangle around y-axis.

OpenStudy (hba):

@Fall12-13

OpenStudy (anonymous):

I'm kinda going rusty on Maths since I don't practice much any more. No idea, bro =[

OpenStudy (hba):

@Fall12-13 I am in a gap-year,so i am having the same problem :(

OpenStudy (hba):

@lgbasallote

OpenStudy (hba):

@Hero

OpenStudy (hba):

@rajathsbhat

hero (hero):

@lgbasallote, all you

OpenStudy (hba):

^ LoL

OpenStudy (lgbasallote):

not my habit to reply in questions already being replied to

OpenStudy (hba):

@lgbasallote Come on dude help me ............

OpenStudy (sirm3d):

let's begin with the formula of the volume of solid of revolution by cylindrical shell: \[\huge V=2\pi \int\limits_{a}^{b}x f(x) dx\]

OpenStudy (sirm3d):



OpenStudy (sirm3d):

x is the distance of the representative strip from the axis of revolution f(x) is the height of the representative strip the horizontal strip sweeps the rectangular region left to right from x= a (the y-axis) to x =b (the x-intercept of the line)